# Phonon effects, impact ionization and power conversion in Mott photovoltaic systems

### v1

@@ IV.1.1 The Hartree shift
We start the analysis of the total exceeding charge $\Delta n_{z}$ in Fig. 6(a) first without considering e-ph interaction. For $\Omega$ up to $\sim 8$ photoexcited electrons from the first two layers $z=1,2$ tunnel to the second half of the heterostructure and leave back a charge deficiency. Such picture is confirmed by the occupation function $N_{z}(\omega)$ in Figs. 4(a)- 5(a). For larger $\Omega$, II is dominant and electrons in the layers $z=1,2$ build-up in doubly occupied states (see Fig. 3(b)-(c)). $\Delta n_{z}$ is strongly influenced by the bias voltage $\Delta V$ which partially fills the UHB of the heterostructure and significantly contributes to the imbalance of $n_{z}$ between the layers (red and blue curves in Fig. 6(a)), especially for those at the interfaces with respect to the ones in the bulk of the correlated region, as already observed in Ref. [31].
Figure 6: (a) Exceeding charge $\Delta n_{z}$ without and with e-ph interaction and (b) Hartree shift $\varepsilon_{\textrm{H},z}$ for the layers $z=1$ and $z=2$, at the e-ph coupling strength $\lambda_{\text{eff}}=0.752$ for different values of the bias voltage $\Delta V$, plotted as functions of the driving frequency $\Omega$. Black dashed vertical lines in (a)-(b) highlight the values taken for the analysis of Figs. 45. Due to the PhI symmetry (see Ref. [31]) $\Delta n_{L+1-z}=-\Delta n_{z}$ and $\varepsilon_{\textrm{H},L+1-z}=-\varepsilon_{\textrm{H},z}$, therefore the curves for $z=3$ and $z=4$ in (a)-(b) are omitted. Default parameters are specified in Table 1. (Here $\Phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
The number of particles per spin per layer $n_{z}$ is influenced also by an another factor which arises from the e-ph interaction. Away from half-filling the so-called Hartree shift, given by the expression derived in Appendix B
$\varepsilon_{\textrm{H},z}=-\frac{2g^{2}}{\omega_{ph}}\Delta n_{z}$ (23)
@@
At $\Delta V=2$ $\varepsilon_{\textrm{H},z=1,2}$, is positive for $\Omega\lesssim 10.5$ [Fig. 6(b)] and the spectral functions of the layers $z=1,2$ are shifted to higher energies, as visible for $\Omega=8$ in Fig. 4(b), while for the layers $z=3,4$ the DOS is moved to lower energies. Overall then, the distance between the peaks of the Hubbard bands increases, as pointed out by the dark green vertical dashed lines in Fig. 4(a). For $\Omega\gtrsim 10.5$ instead, the layers’ DOS has the opposite behavior (not shown).
For $\Delta V=5$ $\varepsilon_{\textrm{H},z=1,2}$ is positive in the whole $\Omega$-range considered. The single layer’s DOS shift is quite large and moves the spectral functions of all layers considerably away from each other, as Fig. 5(a)-(b) shows  [^12] The steep rise in $\Delta n_{z}$ and $\varepsilon_{\textrm{H},z}$ for $\Delta V=5$ with e-ph interaction at $\Omega>15$ (coral curves in Fig. 6) is due to the current flowing in the direction of the bias.. The difference in magnitude between $\varepsilon_{\textrm{H},z,\Delta V=2}$ and $\varepsilon_{\textrm{H},z,\Delta V=5}$, which reflects in a much larger effective potential drop is a direct consequence of the larger $\Delta n_{z}$ for higher bias voltages. $\Delta n_{z}$ is, in turn, strongly influenced by the spectral changes by acoustic phonons, as evident from the curves without and with e-ph interaction in Fig. 6(a)  [^13] In the self-consistent F-DMFT loop, changes in $\Delta n_{z}$ affect $\varepsilon_{\textrm{H},z}$ and viceversa..
IV.2 Photocurrent-bias characteristic and device efficiency
@@
The photocurrent as function of the bias voltage $\Delta V=\mu_{\text{r}}-\mu_{\text{l}}$ is shown in Fig. 7. An electric power gain is obtained in the bottom right quadrant, substantially enhanced by II (for $\Omega=8,11$) with respect to the case of lower driving frequencies ($\Omega=5$). On the other hand, e-ph interaction affects only slightly the curves at small values of $\lambda_{\textrm{eff}}$ (Fig. 7(a)), while at higher $\lambda_{\textrm{eff}}$ the power gain area in the II regime is clearly reduced  [^14] Simulations at $\Omega=5,8$ for the values of $\lambda_{\textrm{eff}}$ in Fig. 7(b) are characterized by large values of $\Delta n_{z}$. This triggers an uncontrolled increase of $\varepsilon_{\textrm{H}}$ in the self-consistent F-DMFT loop and, as consequence, simulations don’t converge. For the same reason, the curves with the e-ph interaction at $\Omega=11$ show, for $\Delta V\gtrsim 6$, an increase of the associated error $\sigma_{j}$., see Fig. 7(b).
@@
Figure 7: (a) Time-averaged steady state current $j$ plotted as function of the bias voltage $\Delta V=\mu_{\text{r}}-\mu_{\text{l}}$ for (a) different values of the driving frequency $\Omega$ without and with e-ph interaction at $\lambda_{\textrm{eff}}=0.212$ and (b) different e-ph coupling strengths $\lambda_{\textrm{eff}}$ at $\Omega=11$. The orange curve in (a) denotes the background current [25, 30, 31] obtained without external driving and without e-ph interaction. Dark green curve in (b) without e-ph interaction is replotted from (a) for comparison. Error bars corresponding to $\sigma_{j}$ are not shown in (a) and only for some values of $\lambda_{\textrm{eff}}$ in (b), for better visualization of the data. Default parameters are specified in Table 1. (Here $\phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
We want to study now the efficiency $\eta$ of the device, starting from the analysis of the time-averaged absorbed power $\overline{P}_{\text{abs}}$ introduced in Sec. III.2. Plotted as a function of $\Omega$, this quantity shows a peak at $\Omega\approx 11$ (see Fig. 8(a)), similarly to the photocurrent $|j|$ [Fig. 3(a)] and is only weakly affected by the increasing bias voltage. E-ph interaction enhances $\overline{P}_{\text{abs}}$ at small $\Omega$ and suppresses it around the peak for large $\Delta V$. Due to the presence of a finite spectral weight within the gap  [^15] A residual DOS in the gap region of the correlated layers is due to the hybridization to the leads. a nonzero contribution of $\overline{P}_{\text{abs}}$ is present also for $\Omega\lesssim\Delta_{\text{g}}$. On the other hand, the time-averaged output power in Fig. 8(b) is computed as $\overline{P}_{\text{out}}\equiv(|j|-|j|_{\text{back}})\Delta V$, where the fictious contribution from the background current $|j|_{\text{back}}$ (see orange curve in Fig. 7(a)) has been removed.
@@
Figure 8: (a) Time-averaged absorbed power $\overline{P}_{\text{abs}}$ and (b) time-averaged output power $\overline{P}_{\text{out}}$ with corresponding error bars, plotted as function of the driving frequency $\Omega$ for different values of the bias voltage $\Delta V=\mu_{\text{r}}-\mu_{\text{l}}$, without and with e-ph interaction at $\lambda_{\textrm{eff}}=0.752$. Default parameters are specified in Table 1. (Here $\Phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
@@
$\eta\equiv\frac{\overline{P}_{\text{out}}}{\overline{P}_{\text{abs}}}$ (24)
provides an assessment for the power conversion quality of the photovoltaic device. $\eta$ shows a peak for $\Omega\approx 8$, as one can see in Fig. 9  [^16] The uncertainties displayed in Fig. 9 are given by $\sigma_{\eta}=\frac{(\sigma_{j}+\sigma_{j_{\text{back}}})\Delta V}{\overline{P}_{\text{abs}}}$., confirming once more the importance of II for an efficient power conversion, in line with the results of Fig. 7(a) (blue and dark green curves). As for the photocurrent $|j|$, $\eta$ is only weakly affected by the e-ph interaction at low $\Delta V$, in contrast to higher $\Delta V$. Even in this case, the role of II as key physical process enhancing power conversion seems to be supported by the peak at $\Omega\approx 10$ displayed in Fig. 9 (dashed blue curve). The finite contribution of $\eta$ for $\Omega\lesssim\Delta_{\text{g}}$ is due to the nonzero $\overline{P}_{\text{abs}}$ and $\overline{P}_{\text{out}}$ at the same frequencies, as already mentioned.
[…]
For these reasons $\eta$ in Eq. (24) should not be compared with the efficiency of a physical device, but should be rather used to qualitatively investigate promising regions of the spectrum for an efficient power conversion.
Figure 9: Efficiency $\eta$ plotted as function of the driving frequency $\Omega$ for different values of the bias voltage $\Delta V=\mu_{\text{r}}-\mu_{\text{l}}$, without and with e-ph interaction at $\lambda_{\text{eff}}=0.752$. Error bars corresponding to $\sigma_{\eta}$ are shown only for some values of $\Delta V$, for better visualization of the data. Default parameters are specified in Table 1. (Here $\Phi=2$, $\Delta_{\text{g}}\approx 3$ and $W_{\text{b}}\approx 8.5$.)
V Conclusions
In this paper we consider a simplified Mott photovoltaic setup which consists of $\text{L}=4$ strongly-correlated layers under the illumination of a monochromatic, peri- odic electromagnetic field with driving frequency $\Omega$, connected to wideband metallic leads with narrowband layers in between. A (Floquet) steady-state photocurrent flows from the left to the right lead against the potential energy originated by the bias voltage applied to the leads’ electrodes. We treat electronic correlations within real-space Floquet dynamical mean-field theory with an accurate steady-state impurity solver.
@@ V Conclusions
The results obtained in this paper are valid for external drivings as ultrashort pulsed lasers [48], with field intensities several orders of magnitude larger than those typical of the sunlight.
In this work our main interest is to address the effects of electron-electron and electron-phonon interaction on transport and spectral properties of the chosen setup. For this reason, other important physical effects are missing in the present study. Future investigations which aim to quantitatively assess the efficiency of Mott photovoltaic devices should also consider those aspects. One example is the conversion of thermal energy into chemical energy, which leads to an additional chemical current [80]. Another one is the effect of the electron-hole pairs recombination due to the Auger effect [11], which manifests in a recombination current and reduces the output power. Also the long-range Coulomb interaction between doublons and holons has to be considered, as well as excitonic effects, which would be important to properly address the dynamics of photoexcited electron-hole pairs and might be described via self-consistent nonequilibrium GW+EDMFT schemes [84, 85]. Moreover, charge reconstruction effects [86, 87, 88] are especially important for Mott photovoltaics, due to the polar interfaces present in the central region, the charge distribution of which needs to be properly considered. In addition to that, Refs. [89, 90] showed that an essential parameter for high-quality Mott photovoltaic setups is the defects concentration: one might then include impurity scattering at the level of the self-consistent Born approximation in a F-DMFT framework with low additional computational effort [51]. Finally, magnetic effects and multi-orbital physics considerably influence the layers’ density of states [24] and change the transport properties, as already shown in Refs. [38, 24].
Acknowledgements.
@@ Appendix A Real-space Floquet DMFT
In this Appendix we summarize the main steps of the self-consistent real-space F-DMFT [57, 59, 47, 22, 33, 36, 39, 40] scheme to obtain the electron and e-ph SE. In real-space F-DMFT one takes into account the electron and e-ph SE as spatially local, i.e. $\underline{{\bf\Sigma}}_{zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{\bf{\Sigma}}_{z}(\omega)\delta_{zz^{\prime}}$, $\underline{{\bf\Sigma}}_{\textrm{e-ph},zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{{\bf\Sigma}}_{\textrm{e-ph},z}(\omega)\delta_{zz^{\prime}}$. Within this approximation, one solves for each correlated layer $z$ a (nonequilibrium) quantum impurity model with Hubbard interaction $U_{z}$ and onsite energy $\varepsilon_{z}$, with a bath hybridization function $\underline{\bm{\Delta}}_{z}(\omega)$ which has to be determined self-consistently.
@@
  • (i) Initial guess for the electron SE $\underline{\bf{\Sigma}}_{z}(\omega)$,
  • (ii) Computation of the local electron GF via the recursive Green’s function method [22, 33, 39, 91, 92, 31], i.e. $\underline{\bf G}_{\text{loc},zz}(\omega)=\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\ \rho(\epsilon,\overline{\epsilon})\underline{{\bf G}}_{zz}(\omega;\epsilon,\overline{\epsilon})$,
  • (iii) Evaluation of the e-ph SE $\underline{{\bf\Sigma}}_{\textrm{e-ph},z}(\omega)$ via Eqs. (11) and (23),
  • (iv) Mapping of the problem onto a single impurity plus bath, with hybridization function $\underline{\bm{\Delta}}_{z}(\omega)=\underline{\bm{g}}^{-1}_{0,z,\text{site}}(\omega)-\underline{\bm{G}}^{-1}_{\text{loc},zz}(\omega)-\underline{\bm{\Sigma}}_{z}(\omega)$, in which $\underline{\bm{g}}^{-1}_{0,z,\text{site}}(\omega)$ is defined as in Eq. (LABEL:eq:inv_non-int_lat_GF_comps) with $\varepsilon_{mn}(\epsilon,\overline{\epsilon})=0$ and $t_{zz^{\prime}}=0$,
  • (v) Solution of the nonequilibrium many-body impurity problem, which provides the new $\underline{\bm{\Sigma}}_{z}(\omega)$,
@@
In step (iv) we consider only the $(0,0)$-Floquet matrix element of all the quantities in the expression for $\underline{\bm{\Delta}}_{z}(\omega)$, because we adopt the Floquet-diagonal self-energy approximation (FDSA) [25, 30, 31] for $\underline{\bf{\Sigma}}_{z}(\omega)$ and $\underline{{\bf\Sigma}}_{\textrm{e-ph},z}(\omega)$, whereby the nondiagonal Floquet indices of such quantities are neglected. The other diagonal components of the SE are obtained by making use of the property $\underline{\Sigma}_{mm}(\omega)=\underline{\Sigma}_{00}(\omega+m\Omega)$. Thanks to the PhI symmetry of the system (see Ref. [31]), we solve the many-body problem in step (v) only for half of the correlated layers, via the auxiliary master equation approach (AMEA) [93, 94, 95, 96, 97, 50, 98, 99].

### v2

@@ 4.1.1 The Hartree shift
We start the analysis of the total exceeding charge $\Delta n_{z}$ in Fig. 6(a) first without considering e-ph interaction. For $\Omega$ up to $\sim 8$ photoexcited electrons from the first two layers $z=1,2$ tunnel to the second half of the heterostructure and leave back a charge deficiency. Such picture is confirmed by the occupation function $N_{z}(\omega)$ in Figs. 4(a)- 5(a). For larger $\Omega$, II is dominant and electrons in the layers $z=1,2$ build-up in doubly occupied states (see Fig. 3(b)-(c)). $\Delta n_{z}$ is strongly influenced by the bias voltage $\Delta V$ which partially fills the UHB of the heterostructure and significantly contributes to the imbalance of $n_{z}$ between the layers (red and blue curves in Fig. 6(a)), especially for those at the interfaces with respect to the ones in the bulk of the correlated region.
Figure 6: (a) Exceeding charge $\Delta n_{z}$ without and with e-ph interaction and (b) Hartree shift $\varepsilon_{\mathrm{H},z}$ for the layers $z=1$ and $z=2$, at the e-ph coupling strength $\lambda_{\mathrm{eff}}=0.752$ for different values of the bias voltage $\Delta V$, plotted as functions of the driving frequency $\Omega$. Black dashed vertical lines in (a)-(b) highlight the values taken for the analysis of Figs. 45. Due to the PhI symmetry [31] $\Delta n_{L+1-z}=-\Delta n_{z}$ and $\varepsilon_{\mathrm{H},L+1-z}=-\varepsilon_{\mathrm{H},z}$, therefore the curves for $z=3$ and $z=4$ in (a)-(b) are omitted. Default parameters are specified in Table 1. (Here $\Phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
The number of particles per spin per layer $n_{z}$ is influenced also by an another factor which arises from the e-ph interaction. Away from half-filling the so-called Hartree shift, given by the expression derived in B
$\varepsilon_{\mathrm{H},z}=-\frac{2g^{2}}{\omega_{ph}}\Delta n_{z}$ (23)
@@
At $\Delta V=2$ $\varepsilon_{\mathrm{H},z=1,2}$, is positive for $\Omega\lesssim 10.5$ [Fig. 6(b)] and the spectral functions of the layers $z=1,2$ are shifted to higher energies, as visible for $\Omega=8$ in Fig. 4(b), while for the layers $z=3,4$ the DOS is moved to lower energies. Overall then, the distance between the peaks of the Hubbard bands increases, as pointed out by the dark green vertical dashed lines in Fig. 4(a). For $\Omega\gtrsim 10.5$ instead, the layers’ DOS has the opposite behavior (not shown).
For $\Delta V=5$ $\varepsilon_{\mathrm{H},z=1,2}$ is positive in the whole $\Omega$-range considered. The single layer’s DOS shift is quite large and moves the spectral functions of all layers considerably away from each other, as Fig. 5(a)-(b) shows  [^12] The steep rise in $\Delta n_{z}$ and $\varepsilon_{\mathrm{H},z}$ for $\Delta V=5$ with e-ph interaction at $\Omega>15$ (coral curves in Fig. 6) is due to the current flowing in the direction of the bias.. The difference in magnitude between $\varepsilon_{\mathrm{H},z,\Delta V=2}$ and $\varepsilon_{\mathrm{H},z,\Delta V=5}$, which reflects in a much larger effective potential drop is a direct consequence of the larger $\Delta n_{z}$ for higher bias voltages. $\Delta n_{z}$ is, in turn, strongly influenced by the spectral changes by acoustic phonons, as evident from the curves without and with e-ph interaction in Fig. 6(a)  [^13] In the self-consistent F-DMFT loop, changes in $\Delta n_{z}$ affect $\varepsilon_{\mathrm{H},z}$ and viceversa..
4.2 Photocurrent-bias characteristic and device efficiency
@@
The photocurrent as function of the bias voltage $\Delta V=\mu_{\mathrm{r}}-\mu_{\mathrm{l}}$ is shown in Fig. 7. An electric power gain is obtained in the bottom right quadrant, substantially enhanced by II (for $\Omega=8,11$) with respect to the case of lower driving frequencies ($\Omega=5$). On the other hand, e-ph interaction affects only slightly the curves at small values of $\lambda_{\mathrm{eff}}$ (Fig. 7(a)), while at higher $\lambda_{\mathrm{eff}}$ the power gain area in the II regime is clearly reduced  [^14] Simulations at $\Omega=5,8$ for the values of $\lambda_{\mathrm{eff}}$ in Fig. 7(b) are characterized by large values of $\Delta n_{z}$. This triggers an uncontrolled increase of $\varepsilon_{\mathrm{H}}$ in the self-consistent F-DMFT loop and, as consequence, simulations don’t converge. For the same reason, the curves with the e-ph interaction at $\Omega=11$ show, for $\Delta V\gtrsim 6$, an increase of the associated error $\sigma_{j}$., see Fig. 7(b).
@@
Figure 7: (a) Time-averaged steady state current $j$ plotted as function of the bias voltage $\Delta V=\mu_{\mathrm{r}}-\mu_{\mathrm{l}}$ for (a) different values of the driving frequency $\Omega$ without and with e-ph interaction at $\lambda_{\mathrm{eff}}=0.212$ and (b) different e-ph coupling strengths $\lambda_{\mathrm{eff}}$ at $\Omega=11$. The orange curve in (a) denotes the background current [25, 30, 31] obtained without external driving and without e-ph interaction. Dark green curve in (b) without e-ph interaction is replotted from (a) for comparison. Error bars corresponding to $\sigma_{j}$ are not shown in (a) and only for some values of $\lambda_{\mathrm{eff}}$ in (b), for better visualization of the data. Default parameters are specified in Table 1. (Here $\phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
@@
We want to study now the efficiency $\eta$ of the device, starting from the analysis of the time-averaged absorbed power $\overline{P}_{\mathrm{abs}}$ introduced in Sec. 3.2. Plotted as a function of $\Omega$, this quantity shows a peak at $\Omega\approx 11$ (see Fig. 8(a)), similarly to the photocurrent $|j|$ [Fig. 3(a)] and is only weakly affected by the increasing bias voltage. E-ph interaction enhances $\overline{P}_{\mathrm{abs}}$ at small $\Omega$ and suppresses it around the peak for large $\Delta V$. Due to the presence of a finite spectral weight within the gap  [^15] A residual DOS in the gap region of the correlated layers is due to the hybridization to the leads. a nonzero contribution of $\overline{P}_{\mathrm{abs}}$ is present also for $\Omega\lesssim\Delta_{\mathrm{g}}$. On the other hand, the time-averaged output power in Fig. 8(b) is computed as $\overline{P}_{\mathrm{out}}\equiv(|j|-|j|_{\mathrm{back}})\Delta V$, where the fictious contribution from the background current $|j|_{\mathrm{back}}$ (see orange curve in Fig. 7(a)) has been removed.
@@
Figure 8: (a) Time-averaged absorbed power $\overline{P}_{\mathrm{abs}}$ and (b) time-averaged output power $\overline{P}_{\mathrm{out}}$ with corresponding error bars, plotted as function of the driving frequency $\Omega$ for different values of the bias voltage $\Delta V=\mu_{\mathrm{r}}-\mu_{\mathrm{l}}$, without and with e-ph interaction at $\lambda_{\mathrm{eff}}=0.752$. Default parameters are specified in Table 1. (Here $\Phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
@@
$\eta\equiv\frac{\overline{P}_{\mathrm{out}}}{\overline{P}_{\mathrm{abs}}}$ (24)
provides an assessment for the power conversion quality of the photovoltaic device. $\eta$ shows a peak for $\Omega\approx 8$, as one can see in Fig. 9  [^16] The uncertainties displayed in Fig. 9 are given by $\sigma_{\eta}=\frac{(\sigma_{j}+\sigma_{j_{\mathrm{back}}})\Delta V}{\overline{P}_{\mathrm{abs}}}$., confirming once more the importance of II for an efficient power conversion, in line with the results of Fig. 7(a) (blue and dark green curves). As for the photocurrent $|j|$, $\eta$ is only weakly affected by the e-ph interaction at low $\Delta V$, in contrast to higher $\Delta V$. Even in this case, the role of II as key physical process enhancing power conversion seems to be supported by the peak at $\Omega\approx 10$ displayed in Fig. 9 (dashed blue curve). The finite contribution of $\eta$ for $\Omega\lesssim\Delta_{\mathrm{g}}$ is due to the nonzero $\overline{P}_{\mathrm{abs}}$ and $\overline{P}_{\mathrm{out}}$ at the same frequencies, as already mentioned.
One should consider that the expression 24 for $\eta$ is valid in the ideal case in which the electromagnetic radiation is completely absorbed by the the device region under illumination [64, 11, 65]. In the present situation $\overline{P}_{\mathrm{abs}}$ is however some orders of magnitude smaller than $\overline{P}_{\mathrm{in}}$, the incoming power per unit surface from a perpendicular electromagnetic radiation, given by the Poynting vector $\overline{P}_{\mathrm{in}}=\frac{1}{2}cn_{\mathrm{refr}}\varepsilon_{0}|E_{0}|^{2}$  [^17] Here, $c$ is the velocity of light, $n_{\mathrm{refr}}$ the refraction index of the medium and $\varepsilon_{0}$ the vacuum permittivity.. This occurs because $\overline{P}_{\mathrm{in}}$ penetrates the central correlated region and is absorbed only over a large number of layers  [^18] From Ref. [43], the absorption coefficient $\alpha$ in LaVO3/SrTiO3 gives a penetration depth of $\sim 10^{4}a$, with $a\sim 10\mathrm{\AA}$ the lattice spacing.. In addition, Eq. 24 holds for a monochromatic driving and therefore, in the case of a solar cell, numerator and denominator should be replaced by a weighted average over the solar spectrum.
For these reasons $\eta$ in Eq. 24 should not be compared with the efficiency of a physical device, but should be rather used to qualitatively investigate promising regions of the spectrum for an efficient power conversion.
Figure 9: Efficiency $\eta$ plotted as function of the driving frequency $\Omega$ for different values of the bias voltage $\Delta V=\mu_{\mathrm{r}}-\mu_{\mathrm{l}}$, without and with e-ph interaction at $\lambda_{\mathrm{eff}}=0.752$. Error bars corresponding to $\sigma_{\eta}$ are shown only for some values of $\Delta V$, for better visualization of the data. Default parameters are specified in Table 1. (Here $\Phi=2$, $\Delta_{\mathrm{g}}\approx 3$ and $W_{\mathrm{b}}\approx 8.5$.)
5 Conclusions
In this paper we consider a simplified Mott photovoltaic setup which consists of $\mathrm{L}=4$ strongly-correlated layers under the illumination of a monochromatic, periodic electromagnetic field with driving frequency $\Omega$, connected to wideband metallic leads with narrowband layers in between. A (Floquet) steady-state photocurrent flows from the left to the right lead against the potential energy originated by the bias voltage applied to the leads’ electrodes. We treat electronic correlations within real-space Floquet dynamical mean-field theory with an accurate steady-state impurity solver.
@@ 5 Conclusions
The results obtained in this paper are valid for external drivings as ultrashort pulsed lasers [46], with field intensities several orders of magnitude larger than those typical of the sunlight. In this work our main interest is to address the effects of electron-electron and electron-phonon interaction on transport and spectral properties of the chosen setup. For this reason, other important physical effects are missing in the present study. Future investigations which aim to quantitatively assess the efficiency of Mott photovoltaic devices should also consider those aspects. One example is the conversion of thermal energy into chemical energy, which leads to an additional chemical current [64]. Another one is the effect of the electron-hole pairs recombination due to the Auger effect [11], which manifests in a recombination current and reduces the output power. Also the long-range Coulomb interaction between doublons and holons has to be considered, as well as excitonic effects, which would be important to properly address the dynamics of photoexcited electron-hole pairs and might be described via self-consistent nonequilibrium GW+EDMFT schemes [66, 67]. Moreover, charge reconstruction effects [68, 69, 70] are especially important for Mott photovoltaics, due to the polar interfaces present in the central region, the charge distribution of which needs to be properly considered. In addition to that, Refs. [71, 72] showed that an essential parameter for high-quality Mott photovoltaic setups is the defects concentration: one might then include impurity scattering at the level of the self-consistent Born approximation in a F-DMFT framework with low additional computational effort [49]. Finally, magnetic effects and multi-orbital physics considerably influence the layers’ density of states [24] and change the transport properties, as already shown in Refs. [38, 24].
Data availability statement
The data that support the findings of this study will be openly available at the following URL/DOI: https://repository.tugraz.at/uploads/e6zzj-93j27.
@@ Appendix A Real-space Floquet DMFT
In this Appendix we summarize the main steps of the self-consistent real-space F-DMFT [55, 57, 45, 22, 33, 36, 39, 40] scheme to obtain the electron and e-ph SE. In real-space F-DMFT one takes into account the electron and e-ph SE as spatially local, i.e. $\underline{{\mathbf{\Sigma}}}_{zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{\mathbf{\Sigma}}_{z}(\omega)\delta_{zz^{\prime}}$, $\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},zz^{\prime}}(\omega;\epsilon,\overline{\epsilon})\approx\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},z}(\omega)\delta_{zz^{\prime}}$. Within this approximation, one solves for each correlated layer $z$ a (nonequilibrium) quantum impurity model with Hubbard interaction $U_{z}$ and onsite energy $\varepsilon_{z}$, with a bath hybridization function $\underline{\underline{\underline{\Delta}}}_{z}(\omega)$ which has to be determined self-consistently.
@@
  • (i) Initial guess for the electron SE $\underline{\mathbf{\Sigma}}_{z}(\omega)$,
  • (ii) Computation of the local electron GF via the recursive Green’s function method [22, 33, 39, 73, 74, 31], i.e. $\underline{\mathbf{G}}_{\mathrm{loc},zz}(\omega)=\int\mathop{}\!\mathrm{d}\epsilon\int\mathop{}\!\mathrm{d}\overline{\epsilon}\ \rho(\epsilon,\overline{\epsilon})\underline{{\mathbf{G}}}_{zz}(\omega;\epsilon,\overline{\epsilon})$,
  • (iii) Evaluation of the e-ph SE $\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},z}(\omega)$ via Eqs. 11 and 23,
  • (iv) Mapping of the problem onto a single impurity plus bath, with hybridization function $\underline{\underline{\underline{\Delta}}}_{z}(\omega)=\underline{\underline{\underline{g}}}^{-1}_{0,z,\mathrm{site}}(\omega)-\underline{\underline{\underline{G}}}^{-1}_{\mathrm{loc},zz}(\omega)-\underline{\underline{\underline{\Sigma}}}_{z}(\omega)$, in which $\underline{\underline{\underline{g}}}^{-1}_{0,z,\mathrm{site}}(\omega)$ is defined as in Eq. 3.1 with $\varepsilon_{mn}(\epsilon,\overline{\epsilon})=0$ and $t_{zz^{\prime}}=0$,
  • (v) Solution of the nonequilibrium many-body impurity problem, which provides the new $\underline{\underline{\underline{\Sigma}}}_{z}(\omega)$,
@@
In step (iv) we consider only the $(0,0)$-Floquet matrix element of all the quantities in the expression for $\underline{\underline{\underline{\Delta}}}_{z}(\omega)$, because we adopt the Floquet-diagonal self-energy approximation (FDSA) [25, 30, 31] for $\underline{\mathbf{\Sigma}}_{z}(\omega)$ and $\underline{{\mathbf{\Sigma}}}_{\textrm{e-ph},z}(\omega)$, whereby the nondiagonal Floquet indices of such quantities are neglected. The other diagonal components of the SE are obtained by making use of the property $\underline{\Sigma}_{mm}(\omega)=\underline{\Sigma}_{00}(\omega+m\Omega)$. Thanks to the PhI symmetry of the system [31], we solve the many-body problem in step (v) only for half of the correlated layers, via the auxiliary master equation approach (AMEA) [75, 76, 77, 78, 79, 48, 80, 81].